Find the Volume of the Cylinder. Round to the nearest tenth.
(1.5-inches height (4.5-inches radius)
help me find the answer.

Answer by checkley75(3666)

VOL OF TOP
A=PIR^2
A=3.14*4.5^2
A=3.14*20.25
A=63.585 AREA OF THE TOP
VOL=AREA OF TOP*HEIGHT
63.585*1.5=95.4 CUBIC INCHES FOR THE VOLUME